OpenStudy (anonymous):

Multiply the expressions, and simplify. (8x3y-3) · (3x3y-3)

OpenStudy (lilg132):

actually i'll let thomas do this one =)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know how to distribute?

OpenStudy (anonymous):

some what yeah

OpenStudy (anonymous):

this is just a distribution problem but it will take a little longer to compute

OpenStudy (anonymous):

okay help me to start it and I should be able to figure it out I know 8*3 =24

OpenStudy (anonymous):

do you add like exponents

OpenStudy (anonymous):

sure your first term will be (8x3y * 3x3y)

OpenStudy (anonymous):

let me expand it out for you and then you can try to see the patteren

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[(8x3y-3) * (3x3y-3) = (8x3y*3x3y) + (8x3y*-3) + (-3*3x3y) +(-3*-3)\]

OpenStudy (anonymous):

did my eqaution go off the screen for you?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

haha sorry the part that is cut off is +(-3 * -3)

OpenStudy (anonymous):

all I did was FOIL it out. Now you have to go about simplying it.

OpenStudy (anonymous):

yeah but what about the exponents the problem was (8x^3y^-3) · (3x^3y^-3) sorry didn't see that when I posted it

OpenStudy (anonymous):

don't you add the exponents when multiplying the main part of the problem

OpenStudy (anonymous):

ah no problem you can slip them into the problem when you mutiply two exponents with the same base you can add the exponents for example \[x^3 * x^5 = x^8\]

OpenStudy (anonymous):

okay so then it is 24x^6y^-6 correct

OpenStudy (anonymous):

yes!!!!

OpenStudy (anonymous):

well done! that one can get messy

OpenStudy (anonymous):

okay thought it was right but wasn't sure thanks again

OpenStudy (anonymous):

yeah especially when it is posted wrong

OpenStudy (anonymous):

I think I have found a new website for tutoring

OpenStudy (anonymous):

everyone is so helpful